OpenStudy (jen543):

1 math question??! Please help! will fan!

OpenStudy (jen543):

Solve (v+5)^2 - 48= 0, where v is a real number Round your answer to the nearest hundredth If there is more than one solution, separate them with commas

OpenStudy (jen543):

I got the answer 1.92 and -11.92, am I correct?

OpenStudy (mathmale):

Please check your results by substituting them back into the original equation: (v+5)^2 - 48= 0. If you end up with a true equation, then your solutions are correct.

jhonyy9 (jhonyy9):

add to both sides 48 like a first step

OpenStudy (mathmale):

Starting with your v=1.92, I round that up to 2, and substitute that 2 into (v+5)^2 - 48= 0. that comes out to 7^2-48, which is close to zero. Thus, I'd expect your v=1.92 is correct. Check it by substition. Please do the same with your v=-1.92.

OpenStudy (jen543):

Okay thanks!

OpenStudy (mathmale):

:)